If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa and Tom worked together to complete it?
Answer by JimboP1977(311)

After 1 day Sam has done 1/4 of the job, Lisa 1/6 and Tom 1/2
x((1/4)+(1/6)+(1/2)) = 1 where x is the number of days for the job to be completed when all three are working.
x = 1/(((1/4)+(1/6)+(1/2))
x = 12/11 or in other words it takes 1 and 1/11 of a day to do the job.